# CLAMP: Crowdsourcing a LArge-scale in-the-wild haptic dataset with an open-source device for Multimodal robot Perception

### v1

@@ Abstract
Robust robot manipulation in unstructured environments often requires understanding object properties that extend beyond geometry, such as material or compliance—properties that can be challenging to infer using vision alone. Multimodal haptic sensing provides a promising avenue for inferring such properties, yet progress has been constrained by the lack of large, diverse, and realistic haptic datasets. In this work, we introduce the CLAMP device, a low-cost (<\$200) sensorized reacher-grabber designed to collect large-scale, in-the-wild multimodal haptic data from non-expert users in everyday settings. We deployed 16 CLAMP devices to 41 participants, resulting in the CLAMP dataset, the largest open-source multimodal haptic dataset to date, comprising 12.3 million datapoints across 5357 household objects. Using this dataset, we train a haptic encoder that can infer material and compliance object properties from multimodal haptic data. We leverage this encoder to create the CLAMP model, a visuo-haptic perception model for material recognition that generalizes to novel objects and three robot embodiments with minimal finetuning. We also demonstrate the effectiveness of our model in three real-world robot manipulation tasks: sorting recyclable and non-recyclable waste, retrieving objects from a cluttered bag, and distinguishing overripe from ripe bananas. Our results show that large-scale, in-the-wild haptic data collection can unlock new capabilities for generalizable robot manipulation. Device assembly guide, dataset, and code can be found here: https://emprise.cs.cornell.edu/clamp/
Figure 1: We present the CLAMP device for collecting multimodal haptic data in the wild. Using data from 16 devices, we build the CLAMP dataset and train models for material & compliance recognition. Our models generalize to different robot embodiments and enable robust robot manipulation in real-world settings.
@@ 1 Introduction
Robots benefit from knowledge of object properties beyond geometry, such as material and compliance, to manipulate objects in real-world scenarios [1, 2, 3] such as sorting waste or grasping deformable objects. However, it is challenging to reliably determine object material or compliance using vision alone [4, 5]. Haptic sensing is informative towards recognizing these object properties, both by itself [6, 7, 8, 9] and coupled with vision [10, 11], but it remains less utilized in robotics than vision or language partly due to the limitations of existing haptic datasets. Current datasets [12, 13, 14, 15] lack one or more of the following attributes:
@@
    Multimodality: Most state-of-the-art haptic datasets capture only contact force [12, 16, 17]. However, haptic sensing is multimodal, including modalities such as temperature [18, 19], force [20, 21], vibration [22, 23], and proprioception [24]. More haptic modalities improve performance of object instance recognition [25] and compliance recognition [15] methods.
@@
    Scalability: Haptic data requires physical interaction, making collection resource-intensive. Collecting haptic data from multiple users can enable large-scale datasets, but requires standardized sensing hardware that is easy to build and use.
@@
    Diversity: Most haptic datasets are collected in controlled environments with a curated list of objects [22, 14]. These datasets do not capture real-world haptic interactions during manipulation. Moreover, these datasets do not capture interactions with heterogeneous household objects, i.e., objects made of multiple materials (e.g. mobile phones, packaged snacks).
@@
    Data from Grasping Interaction: Haptic sensing is action-conditioned, i.e., it depends on the actions an agent takes while interacting with an object. Most datasets focus on non-prehensile interaction with one virtual finger [7, 26, 15]. As a result, their data differ from that generated during real-world manipulation involving grasping.
In this work, we crowdsource a large, in-the-wild haptic dataset for multimodal robot perception. Our key insight is that large-scale multimodal haptic data collected in the wild can help to create perception models that enable robust robot manipulation for different embodiments. Our contributions are as follows:
@@
    The CLAMP device, a low-cost sensorized reacher-grabber that captures five haptic modalities: active thermal, passive thermal, force, vibration, and proprioceptive sensing. The device can store data, record image and speech annotations, and is designed for long-term deployment with non-expert users: it is lightweight (1.3 lbs) and contains a display with a GUI that allows users to collect data easily with clear instructions.
@@
    The CLAMP dataset — the largest open-source multimodal haptic dataset in the robotics literature — containing 12.3 million data samples from 5357 in-situ household objects gathered by 41 non-expert users with a total of 16 CLAMP devices. The CLAMP dataset contains data from a wide range of household objects with diversity in grasp orientation, speed, and force.
@@
    The CLAMP model, a visuo-haptic model for material classification that generalizes to novel objects and data from multiple robots and gripper types with minimal finetuning. Additionally, we show that the haptic encoder within the CLAMP model can be transferred to classify objects based on compliance with no finetuning.
@@
    Real robot experiments that demonstrate a finetuned CLAMP model with a 7-DoF Franka Panda in three real-world manipulation tasks: sorting recyclable and non-recyclable waste, where the robot reasons about object material, retrieving objects from a cluttered bag, where the robot reasons about object material under occlusion and visual ambiguity, and separating ripe from overripe bananas, where the robot reasons about object compliance.
@@ 2 Related Work
Data acquisition in haptics. Prior research in object instance recognition and material recognition has collected haptic data from robots [7, 14, 17, 28, 29] or human experts [12, 14, 26], however this approach is constrained by the availability of robots and human experts. While haptic sensing rigs [30, 31, 16] provide a pathway to scalable data collection, existing rigs are not optimized for size [30], user-friendliness [31], or cost [16], and also lack streamlined workflows for data collection. In contrast, the CLAMP device is designed for deployment in homes and used by non-experts.
@@
Haptic datasets. State-of-the-art haptic datasets [12, 13, 14, 15, 7, 32, 33, 34, 25, 17, 28] contain fewer data samples than those in vision [35], language [36] or robot actions [37]. Previously, Yang et al. [12] created the largest haptic dataset by number of objects, consisting of 3971 objects, while Cheng et al. [13] aggregated the largest dataset by data points with 100k object touches. In comparison, the CLAMP dataset contains 5357 household objects and 12.3 million multimodal data samples, representing a sizable increase in data scale.
@@
Most datasets are collected using a pre-determined set of materials [18, 38, 19] or objects [15, 33, 17], typically by robots or human experts who ensure uniform interaction and consistent contact duration [15, 12, 33, 17, 39, 7]. In comparison, our data collection places no restrictions on objects or materials and allows users to grasp items freely, enabling diverse in-situ haptic data. Finally, prior haptic datasets capture a subset of haptic modalities, such as force [20, 40, 12, 26, 16], vibration [23], and thermal [18, 41] sensing, while some include a combination of the above [15, 7, 32, 33, 34]. The CLAMP dataset captures 5 haptic modalities along with vision and structured language.
Object property recognition. Prior works use vision, [4, 42], haptics [26, 7], or both [10] to reason about object properties such as object instance, material and compliance. Works in computer vision note that material recognition is challenging using only vision [43, 4]; recent work on grounding VLMs [42, 44] improves object property recognition but highlights multimodal reasoning as a key next step. Prior work on identifying object instances [10], haptic adjectives [7], and properties [21, 15] uses haptic data from controlled settings. In contrast, our visuo-haptic model is trained on in-situ, real-world multimodal haptic data. Yang et al. [12] collect high-resolution force data in-situ for material recognition but do not incorporate other haptic modalities.
@@
Haptic data is action-conditioned, which makes transferring haptic models across robots challenging. Yu et al. [26] show that their tactile model, trained on human data, transfers to robot data for avocado ripeness classification, but do not evaluate generalization across objects, embodiments, or reasoning tasks. Tatiya et al. [45] learn a common latent space to transfer haptic knowledge between robots but validate their methods on simulated robots only. In comparison, we transfer our visuo-haptic model from human-collected data to three robot embodiments and show improved real-world robot manipulation in different scenarios.
@@ 3 The CLAMP Device
Figure 2: Device overview: The CLAMP device features a modified reacher-grabber equipped with sensors that capture five haptic modalities. Designed to be easy to build, easy to carry, easy to use, and easy to scale, our device enables non-expert users to collect haptic data in-the-wild.
Sensing. The CLAMP device features a sensorized reacher-grabber that captures a total of five modalities upon contact with an object. The left sensor bed includes a passive unheated thermal sensor that records the ambient temperature of the object, and an active thermal sensor which is heated to 55°C and provides an indication of the heat capacity of the object. The suction cup on the right end-effector contains a contact microphone that measures vibrations, a force-sensing resistor that measures normal force, and a 6DoF IMU that tracks the motion of the right end-effector. These sensors are embedded in 3D printed sensor beds as shown in Figure 2. The device also includes a second IMU mounted on the device body frame which helps to isolate the motion of the end-effector.
Compute and storage. Our device integrates a Raspberry Pi Pico microcontroller and a Raspberry Pi Zero 2W single-board computer (SBC). The Pico is mounted on a custom printed circuit board (PCB) and samples data from the sensors at a rate of 50 Hz for all sensors except the contact microphone, which is sampled at 100Hz. Data is transferred to the SBC via UART. All data is stored on-board and retrieved when the device is returned after deployment.
User Interface. The SBC is connected to an Adafruit 2.2” PiTFT Display. The SBC takes input from four buttons on the PiTFT and powers the display using PyGame. On booting, the computer loads a graphical user interface (GUI) that allows the user to interact with the device and guides them through the data collection process.
Design Principles. We designed the CLAMP device to facilitate large-scale data collection by non-expert users in their homes. We designed the device to be:
@@
    Easy to Build: The CLAMP device can be assembled from scratch within 5 hours (excluding time for 3D printing). We use commercial off-the-shelf and open-source hardware for the CLAMP device, with complete CAD files and assembly instructions provided.
@@
    Easy to Carry: The CLAMP device is lightweight (0.59 kg), making it easy for users to take the device home and collect data with one hand. We achieved this by selecting the lightweight Pi Zero 2W as our onboard computer and using compact end-effectors that keep the center of gravity of the device close to the handle.
@@
    Easy to Use: Users grasp objects with the reacher grabber, receive contact feedback via LEDs, and follow on-screen prompts through the GUI to collect data and annotations. Buttons on the PiTFT allow users to track progress and power down the device.
@@
    Easy to Scale: The total cost of building a CLAMP device is <\$$<\$200$, which is less than that of commercially-available tactile sensors [46, 47, 48]. The sensorized suction cups can be mounted on other robot end-effectors, allowing roboticists to collect haptic data with their own robots.
@@ 4 The CLAMP Dataset
The CLAMP dataset is the largest open-source multimodal haptic dataset to date. The CLAMP dataset consists of 5357 object instances, 25.1k object touches, and 12.3M individual data samples. We gathered this data from 16 CLAMP devices used by 41 people.
Data collection process. Each user is given a short tutorial and a manual on how to use the device. The GUI guides users through three steps of data collection for an object: capturing an image of an object, recording a spoken annotation and collecting haptic data by grasping the object during each trial, for five trials. The GUI then returns to the start screen. At the end of the deployment, the device is returned for maintenance and data backup.
Data Post-Processing and Featurization. We collect data from six haptic sensors during each trial: active and passive thermal sensors, a force sensor, a contact microphone, and two IMUs. After synchronizing all sensor streams, we apply smoothing filters and extract features for model input. For the thermal sensors, the force sensor, and the contact microphone, we use the smoothed readings directly. For the IMU data, we compute the angular velocity of the end-effector IMU relative to the IMU in the gripper body. This results in five primary features. We create four additional features: differences in active thermal, passive thermal, and force readings between consecutive timesteps, and an impedance feature. We create the impedance feature by dividing the force difference feature by angular velocity at each timestep. This ratio reliably indicates contact and shows distinct responses for contact with rigid and compliant objects. We synchronize contacts for each finger by detecting contact onset and loss for the left and right fingers separately. We elaborate on the rule-based approach used to classify contact for each finger, in the appendix. Finally, we pad the features to ensure all features are of the same length.
@@
    One of 16 material labels: {aluminium, brass, cardboard, dry wall, fabric, foam, glass, granite, hard plastic, paper, porcelain, rubber, soft plastic, steel, vegetable matter, wood}. We create this list by starting with 300 material subcategories from the CES Edupack that appear in our dataset, and then group subcategories that are less informative for downstream manipulation. Our material list covers $93\%$ of objects in our dataset.
@@
    Whether the object has different materials on opposing surfaces, i.e., heterogeneous surfaces: {yes, no}. Note the difference between heterogeneous objects (multiple materials overall) and objects with heterogeneous surfaces (two materials on opposite sides, e.g., mobile phones).
@@
To generate these labels, we first transcribe the audio annotation using Whisper [49]. We then provide the audio transcription along with the image annotation for the object to GPT-4o [50], along with more than 12 in-context examples for each label category. We detail the exact prompts used in the appendix. Finally, expert annotators verify the audio transcription and material labels and re-run the annotation pipeline for any corrections made.
Dataset Statistics. The CLAMP dataset contains diversity in objects and in grasping actions. In the appendix, we quantify diversity across three axes: materials, grasping forces, and grasping speeds.

### v2

@@ Abstract
Robust robot manipulation in unstructured environments often requires understanding object properties that extend beyond geometry, such as material or compliance—properties that can be challenging to infer using vision alone. Multimodal haptic sensing provides a promising avenue for inferring such properties, yet progress has been constrained by the lack of large, diverse, and realistic haptic datasets. In this work, we introduce the CLAMP device, a low-cost (<\$200) sensorized reacher-grabber designed to collect large-scale, multimodal, in-the-wild haptic data from non-expert users in everyday settings. We aggregate data from 16 CLAMP devices deployed to 41 users to create the CLAMP dataset, the largest open-source multimodal haptic dataset to date, comprising 12.3 million data points across 5357 household objects. Using this dataset, we train a haptic encoder that can infer object material and compliance from multimodal haptic data. We leverage this encoder to create the CLAMP model, a visuo-haptic perception model for material recognition that generalizes to novel objects and three robot embodiments with minimal finetuning. We also demonstrate the effectiveness of our model in three real-world robot manipulation tasks: sorting recyclable waste from trash, retrieving objects from a cluttered bag, and distinguishing overripe from ripe bananas. Our results show that large-scale, multimodal, in-the-wild haptic data collection can unlock new capabilities for generalizable robot manipulation. Device assembly guide, dataset, and code can be found here: https://emprise.cs.cornell.edu/clamp/
Figure 1: We present the CLAMP device for collecting multimodal, in-the-wild haptic data. Using data from 16 devices, we build the CLAMP dataset and train models for material and compliance recognition. Our models generalize to three different robot embodiments and enable real-world robot manipulation.
@@ 1 Introduction
Robots benefit from knowledge of object properties beyond geometry, such as material and compliance, to manipulate objects in real-world scenarios such as sorting waste or grasping deformable objects [3, 5, 56]. However, it is challenging to reliably determine object material or compliance using vision alone [25, 34]. Haptic sensing helps in recognizing these object properties, both by itself [19, 18, 40, 24] and coupled with vision [30, 37], but it remains less utilized in robotics than vision or language. This is because state-of-the-art haptic datasets [58, 17, 32, 7] lack one or more of the following attributes:
@@
    Multimodality: Most state-of-the-art haptic datasets capture only contact force [58, 28, 16]. However, haptic sensing is inherently multimodal, including temperature [11, 10], force [9, 26], vibration [22, 51], and proprioception [50]. Prior work has shown that using more haptic modalities improve recognition of object properties such as instance [27] and compliance [7].
@@
    Scalability: Haptic data requires physical interaction, making data collection resource-intensive. Collecting haptic data from multiple users can enable large-scale datasets, but requires standardized sensing hardware that is easy to build and use.
@@
    Diversity: Most haptic datasets are collected in controlled environments with a curated list of objects [22, 32]. These datasets do not capture real-world haptic interactions during manipulation. Moreover, these datasets do not capture interactions with heterogeneous household objects, i.e., objects made of multiple materials (e.g. mobile phones, shoes, books with covers).
@@
    Data from Grasping Interaction: Haptic sensing is action-conditioned, i.e., it depends on the actions an agent takes while interacting with an object. Most datasets focus on non-prehensile interaction with one virtual finger [18, 59, 7]. As a result, their data differ from data generated during real-world manipulation involving grasping.
In this work, we crowdsource a large, multimodal, in-the-wild haptic dataset for multimodal robot perception. Our key insight is that large-scale, multimodal, in-the-wild haptic data can help to create perception models that enable real-world robot manipulation for different embodiments. Our contributions are as follows:
@@
    The CLAMP device, a low-cost, sensorized reacher-grabber that captures five haptic modalities: active thermal, passive thermal, force, vibration, and proprioceptive sensing. The device can store data, record image and speech annotations, and is designed for long-term deployment with non-expert users: it is lightweight (0.59 kg), portable, and contains a display with a GUI that allows users to collect data easily with clear instructions.
@@
    The CLAMP dataset, the largest open-source multimodal haptic dataset in the robotics literature — containing 12.3 million data samples from 5357 in-situ household objects gathered by 41 non-expert users with a total of 16 CLAMP devices. The CLAMP dataset contains data from a wide range of household objects with diversity in grasp orientation, speed, and force.
@@
    The CLAMP model, a visuo-haptic model for material classification that generalizes to novel objects and data from multiple robots and gripper types with minimal finetuning. Additionally, the haptic encoder in the CLAMP model can be transferred to classify objects based on compliance with no finetuning.
@@
    Real robot experiments that demonstrate a finetuned CLAMP model with a 7-DoF Franka Panda in three real-world robot manipulation tasks: sorting waste into trash and recycling, where the robot reasons about object material in cluttered scenes, retrieving objects from a cluttered bag, where it reasons about object material under occlusion and visual ambiguity, and separating ripe from overripe bananas, where it reasons about object compliance among objects of the same instance.
@@ 2 Related Work
Data acquisition in haptics. Prior works in object instance recognition and material recognition have collected haptic data from robots [18, 32, 16, 15, 48] or human experts [58, 32, 59], however this approach is constrained by the availability of robots and human experts. While haptic sensing rigs [13, 55, 28] provide a pathway to scalable data collection, existing rigs are not optimized for size [13], user-friendliness [55], or cost [28], and also lack streamlined workflows for data collection. In contrast, the CLAMP device is designed for non-experts to use in their homes.
@@
Haptic datasets. State-of-the-art haptic datasets [58, 17, 32, 7, 18, 33, 12, 54, 27, 16, 15] contain fewer data samples than those in vision [57], language [45] or robot actions [21]. Previously, Yang et al. [58] created the largest haptic dataset by number of objects, consisting of 3971 objects, while Cheng et al. [17] aggregated the largest dataset by data points with 100k object touches. In comparison, the CLAMP dataset contains 5357 household objects and 12.3 million data samples, representing a significant increase in data scale.
@@
Most datasets are collected using a curated set of materials [11, 52, 10] or objects [7, 12, 16], typically by robots or human experts who ensure uniform interaction and consistent contact duration [7, 58, 12, 16, 38, 18]. In comparison, our data collection allows users to collect data from any graspable household object, enabling diverse, in-situ haptic data. Finally, prior datasets capture some haptic modalities among force [9, 8, 58, 59, 28], vibration [51], thermal [11, 4] sensing, or a combination of the above [7, 18, 33, 12, 54]. In comparison, our dataset captures a total of 7 modalities, including 5 haptic modalities along with vision and structured language.
Object property recognition. Real-world robot manipulation can benefit from recognizing object properties in the wild [58]. Recognizing object material from vision alone is challenging without more information [6, 25]. Recent works that ground vision–language models in object properties [29, 1] also highlight multimodal perception as the next step. While many works have used haptic recognition models to identify various object properties [30, 18, 26], these models are trained on data from controlled settings. Yang et al. [58] train a material classification model on unimodal, in-the-wild haptic data, specifically high-resolution force data, collected from non-prehensile actions. In comparison, we train material and compliance classification models on multimodal, in-the-wild haptic data collected from grasping actions, enabling recognition for real-world robot manipulation.
@@
Haptic data is action-conditioned, which makes transferring haptic recognition models across robots challenging. As a result, most work on object property recognition from visuo-haptic sensing [30, 27] pretrain or finetune their models on data from only one robot. Yu et al. [59] show that their tactile model, pretrained on human-collected haptic data, transfers to classification tasks on robot data, but does not show generalization to multiple robots. Tatiya et al. [53] learn a common latent space to transfer haptic knowledge between robots but validate their methods on simulated robots only. In comparison, we show that our visuo-haptic model, pretrained on human-collected data, transfers to three robot embodiments with varying gripper embodiments.
@@ 3 The CLAMP Device
Sensing. The CLAMP device features a sensorized reacher-grabber that captures a total of five modalities upon contact with an object. The left sensor bed includes a passive unheated thermal sensor that records the ambient temperature of the object, and an active thermal sensor heated to 55 °C, whose response over time during contact reflects the object’s heat capacity. The suction cup on the right end-effector houses a contact microphone that measures vibrations, a force-sensing resistor that measures normal force, and a six-axis IMU. These sensors are embedded in 3D printed sensor beds as shown in Figure 2. The device also includes a second IMU, mounted on the device body frame, which helps to isolate the motion of the end-effector.
Compute and storage. Our device integrates all sensors with a Raspberry Pi Pico microcontroller (referred to as Pico) and a Raspberry Pi Zero 2W (referred to as SBC, or single-board computer). The Pico is mounted on a custom printed circuit board (PCB) and samples at a rate of 50 Hz for all sensors except the contact microphone, which is sampled at 100 Hz. Data is transferred to the SBC via a Universal Asynchronous Receiver–Transmitter (UART) interface. All data is stored on-board and retrieved when the device is returned after deployment.
User interface. The SBC is connected to an Adafruit 2.2” PiTFT Display. The SBC takes input from four buttons on the PiTFT and powers the display using PyGame. On booting, the computer loads a graphical user interface (GUI) that allows the user to interact with the device and guides them through the data collection process. Two LEDs, one for each suction cup, provide users with real-time feedback about contact between object surface and each cup, based on sensor readings from the active thermal sensor (right cup) and the force sensor (left cup).
Design. The device’s long handle allows users with limited reach to grasp objects in-situ. We mount the lighter Pico frame along the device length and the heavier SBC enclosure close to the handle, to keep the device ergonomic. As a result, users noted that they held the device in one hand as they repositioned objects with their free hand.
Design principles. We designed the device to be:
@@
    Easy to Build”: The CLAMP device can be assembled from scratch within 5 hours, which is on par with other open-source robotics hardware [61, 49]. The CLAMP device uses readily available sensors sold with datasheets. We also release all CAD files and assembly instructions.
@@
    Easy to Carry”: The CLAMP device is lightweight (0.59 kg), allowing users to take the device home. We achieved this by selecting the lightweight Pi Zero 2W as our onboard computer and using compact end-effectors that keep the center of gravity of the device close to the handle.
@@
    Easy to Use”: Users grasp objects with the reacher-grabber, receive contact feedback via LEDs, and follow on-screen prompts through the GUI to collect data and annotations. Buttons on the PiTFT allow users to track progress and power down the device when not in use.
@@
    Easy to Scale”: One CLAMP device costs $\mathdollar 200$ to build, making it more affordable than existing haptic sensors [35, 36, 60] and sensing suites [20]. The sensorized suction cups can be mounted on robot end-effectors, allowing roboticists to collect haptic data with their own robots.
@@ 4 The CLAMP Dataset
The CLAMP dataset is the largest open-source multimodal haptic dataset to date. The dataset consists of 5357 object instances, 25.1k object touches, and 12.3M individual data samples. We gathered this data from 16 CLAMP devices deployed to 41 users.
Data collection process. Each user is given a short tutorial and a manual on how to use the device. The GUI guides users through three steps of data collection for an object: capturing an image of the object, recording a spoken annotation, and collecting haptic data by grasping the object for up to 10 seconds per trial, for five trials. The GUI then returns to the start screen. At the end of the deployment, the device is returned for maintenance and data backup.
Data post-processing and featurization. We collect data from six haptic sensors during each trial: active and passive thermal sensors, a force sensor, a contact microphone, and two IMUs. After synchronizing all sensor data streams, we apply smoothing filters and extract features from each sensor, resulting in a total of nine features. Refer to Appendix B.1 for more details. We synchronize contacts for each suction cup by detecting contact onset and loss for the left and right cups separately. We elaborate on the rule-based approach used to detect contact onset and release for each cup, in Appendix B.2. Finally, we pad the features to ensure all features are of the same length.
@@
    One of 16 material labels: {aluminium, brass, cardboard, dry wall, fabric, foam, glass, granite, hard plastic, paper, porcelain, rubber, soft plastic, steel, vegetable matter, wood}. We create this list from 300 material subcategories from the CES Edupack [2] database that appear in our dataset, which we group by similarity towards downstream manipulation. Our material list covers $93\%$ of objects in our dataset.
@@
    If the object has different materials on opposing surfaces, i.e., heterogeneous surfaces: {yes, no}. Note the difference between these objects (e.g. mobile phones) and heterogeneous objects (multiple materials overall, e.g. furniture).
@@
To generate these labels, we first transcribe the audio annotation using Whisper [44]. We then provide the audio transcription along with the image annotation for the object to GPT-4o [42], along with more than 12 in-context examples. We detail the exact prompts used in Appendix B.4. Finally, expert annotators verify the audio transcription and material labels and re-run the annotation pipeline for any corrections made.
Dataset statistics. The CLAMP dataset contains diversity in objects and in grasping actions. In the appendix, we quantify diversity across three axes: materials, grasping forces, and grasping speeds. Refer to Appendix B.5 for more details.
Figure 2: Device overview: The CLAMP device features a modified reacher-grabber equipped with sensors that capture five haptic modalities. Our device enables non-expert users to collect haptic data in their homes.